OpenStudy (anonymous):

Solve t = p+3r over 6 for r

OpenStudy (swissgirl):

Solve \( t= \huge\frac{p+3r} {6} \) for r

OpenStudy (swissgirl):

ok so like I said you we have to isolate the r

OpenStudy (swissgirl):

ok now this looks scary but how would we get rid of the 6 at the bottom over there?

OpenStudy (swissgirl):

ok like inorder to "to get rid of" a variable or some number we usually gotta do the opposite. That may have gotten confused but give a wild guess what u think shld be done?

OpenStudy (anonymous):

@Mdias5 ? what exactly did you need help with, and is @swissgirl making sense to you?

OpenStudy (swissgirl):

lol hahahahhaha i didnt make sense lol?!?!?!?

OpenStudy (anonymous):

No, you did. I was just asking them x3

OpenStudy (swissgirl):

Well I solved with him/her 2 other examples

OpenStudy (anonymous):

Oh okay

OpenStudy (swissgirl):

but idk maybe I shld give another hint

OpenStudy (swissgirl):

or maybe I am simply not clear

OpenStudy (allank):

Still there @Mdias5 ? Have you understood @swissgirl ?

OpenStudy (allank):

Hmm...guess you have. She is pretty clear. Maybe in another phrasing, make r the subject of the equation. i.e. rewrite the equation like "r=....."

OpenStudy (swissgirl):

i know what to do here. Like isolating a variable is am important skill so i kinda dont want to give away the answer.

OpenStudy (anonymous):

i understand what to do but i dont know how to do it...if that makes any sense im like the biggest failure at math ever....

OpenStudy (swissgirl):

Well the first step would be to multiply both sides by 6 \( t=\huge \frac{p+3r} {6} \) \(6*t=\huge \frac{p+3r} {6} *6 \)

OpenStudy (anonymous):

The side with r is divided by 6, so you would multiply both sides by 6 to move it

OpenStudy (anonymous):

which then would give you 6t = p+3r

OpenStudy (swissgirl):

Now if you look at the right hand side of the equation you will notice that the 6's will cross each other out and you will be left with: \( \huge 6t=p+3r \)

OpenStudy (swissgirl):

now how can we get rid of that p from the right side of the equation?